OpenStudy (anonymous):

How do I simplify? 4/(1+i) the answer is supposed to be 2-2i ??

OpenStudy (amistre64):

conflagerate the bottom....

OpenStudy (anonymous):

multiply the top and bottom by the conjugate of (1+i) which is (1-i). Simplify.

OpenStudy (amistre64):

\[\frac{4}{1+i} * \frac{1-i}{1-i}\]

OpenStudy (amistre64):

conjugate! thats the word I wanted lol

OpenStudy (anonymous):

4-4i/1+i-i-i2 i2=1 so 4-4i/2 simplify to get 2-2i

OpenStudy (anonymous):

i2=-1

OpenStudy (anonymous):

Awesome, thank you so much for your help!

OpenStudy (anonymous):

welcome from the both of us!